OpenStudy (anonymous):

What is the radius of the circle shown below? x^2+y^2-12x-6y+9=0

OpenStudy (anonymous):

6

OpenStudy (anonymous):

okay we have a formula for this \[r = \sqrt{g^2 + f^2 -c}\] \[r = \sqrt{ 36 + 9 - 9}\] \[r = 6\]

OpenStudy (anonymous):

stephh any problem

OpenStudy (anonymous):

nope, thanks!